# Vertex Representation of Hyperbolic Tensor Networks

### v1

@@ IV Results
We selected the clock and $q$-state Potts models because they exhibit first-order (discontinuous), second-order (continuous), and infinite-order Berezinskii-Kosterlitz-Thouless (BKT) BKT1 ; BKT2 ; BKT3 transitions on the square lattice $(4,4)$. The hyperbolic geometry, however, affects the above results since the Hausdorff dimension of the hyperbolic lattices is infinite. To verify the vertex representation of TN on the hyperbolic lattices, we classify the spin models by phase transition, as stated by the weight representation Ueda1 ; weightpq ; Baxter . We proceed with the following:
@@
Now we check if the vertex representation of TN is correctly constructed. We calculate the phase-transition temperature for the Bethe lattice. The Bethe lattice with the coordination number $r=4$ can be accessed via the set of hyperbolic lattices and is equivalent to $(\infty,4)$ as also examined in Refs. weightpq, ; MFU, ; triangular, . The phase-transition temperature is exactly known Baxter to be $T_{\rm pt}^{(\infty,4)}=2/\ln(2)$. The convergence to the Bethe lattice is common for both the quantum and the classical spin models we have studied in our earlier works by the weight representation only AGquantum ; Mi1 ; Mi2 ; weightp4 .
Figure 8: The temperature dependence of the spontaneous magnetization towards the asymptotic Bethe lattice geometry $(\infty,4)$. The left inset shows the fast asymptotic convergence of the entanglement entropy to the Bethe lattice. The right inset demonstrates the linearity, ${\cal M}^{2}_{p}\propto T$ leading to the mean-field exponent $\beta=\frac{1}{2}$.
@@
Therefore, we show Fig. 9 to display the entanglement entropy in the semi-logarithmic scale to emphasize the differences among the various $(p,4)$ lattices. The decreasing peaks of ${\cal E}_{p}$ completely coincide with the transition temperatures in ${\cal M}_{p}$. In both cases, ${\cal M}_{p}$ and ${\cal E}_{p}$ Equally rapid approach the exact value $T_{\rm pt}^{(\infty,4)}=2/\ln(2)$ on the Bethe lattice.
Figure 9: The entanglement entropy versus temperature in lin-log scale for lattices $(p,4)$ where $4\leq p\leq 50$. The asymptotic convergence of $T_{\rm pt}^{(p,4)}$ to the Bethe lattice is well visible on the sharp maxima of ${\cal E}_{p}$.
Figure 10: Ising model ($q=2$ clock model) calculated on the hyperbolic $(20,4)$ lattice is numerically indistinguishable from the Bethe lattice. The mean-field exponent $\beta=\frac{1}{2}$ can be read-off from the linear decrease of ${\cal M}_{20}^{2}$. The data points (asterisks) shown in the inset required $k>10^{11}$ iteration cycles to converge at $m=2$.
@@ IV.3 Multi-state spin models
Figure 11: The entanglement entropy ${\cal E}_{5}$ vs temperature for the $q$-state clock models on the lattice $(5,4)$. The inset shows the temperature dependence of spontaneous magnetization $M_{5}$ for the $q$-state clock models (lines) that coincides with the results of the weight-type TN, shown by the symbols ($\ast$).
We point out the non-diverging entropy ${\cal E}_{5}$ in the clock model for $q\geq 5$, as plotted in Fig. 11. Recall that ${\cal E}_{4}$ for the $q$-state clock models with $q\geq 2$ on the Euclidean ($4,4$) lattices logarithmically diverges at phase transitions MP ; Ueda ; CentrCharge . On the other hand, phase transitions on the hyperbolic lattices are always non-critical MFU ; triangular since the correlation length $\xi$ does not diverge Iharagi , reaches a maximal sharp peak, and is small ($\xi<1$).
Entanglement entropy at phase transition for $h=b=0$ diverges logarithmically (${\cal E}_{4}\propto\frac{1}{12}\ln k$) on the Euclidean two-dimensional lattices CentrCharge , while ${\cal E}_{p\geq 5}$ is always finite on the hyperbolic lattices $0<{\cal E}_{{p\geq 5}}\ll 1$. The correlations decay exponentially if measured between the bulk and the lattice boundary. However, the correlations on the lattice boundary always decay as a power law and the system is critical on the hyperbolic boundary TNAds1 . Comparing the qualitative similarity between the non-diverging ${\cal E}_{5}$ with the logarithmically diverging ${\cal E}_{4}$ on the square lattice MP , we can conjecture a non-critical BKT-like phase that exists in bulk only. However, no BKT phase transition is present for $p\geq 5$.
@@
Figure 12: The entanglement entropy ${\cal E}_{5}$ vs temperature for the $q$-state Potts models on $(5,4)$ lattice. We emphasize the emergence of the first-order transition for $q\geq 3$ on the hyperbolic lattices. The circles denote the phase-transition temperatures where the curves become discontinuous. Inset shows the equivalence of spontaneous magnetization ${\cal M}_{5}$ between the vertex TNs (lines) and the weight TNs ($\ast$).
@@ IV.4 Free Energy
Figure 13: The temperature dependence of the free energy for the Ising model ($q=2$) on $(p,4)$ lattices when imposing the boundary field $b=0$ and $b=\pm 2$ (at $h=0$). The free energy does not depend on the boundary field $b$ on the square ($4,4$) lattice. The upper inset depicts a strictly linear dependence of the free energy on $b$ and $h$ at $T\to 0$ if studied on the Bethe lattice. The lower inset confirms the high-temperature regime of the thermodynamic entropy $S_{\rm th}\to\ln(q)$. The full and dashed lines correspond to the $q$-state clock and Potts models, respectively.
@@ V Conclusion
Number of the physical spin indices $x_{\cal V}^{\leavevmode\nobreak\ }$, $x_{\cal P}^{\leavevmode\nobreak\ }$, $x_{\cal C}^{\leavevmode\nobreak\ }$ depends on the TN structure which is $p$, $2p$, $p(2p-7)$, respectively, in accord with the right-hand-side of Fig. 15. Such a construction is inevitable for the calculation of $\langle\psi|{\cal H}|\psi\rangle_{(p,4)}$ and other observables on hyperbolic lattices. This way, we access the open questions in the AdS/CFT correspondence ACC0 ; ACC1 .
We commenced the classification of the regular hyperbolic (AdS) lattices viewed from the perspective of condensed matter physics. Our motivation is to expand this study further by evaluating entanglement entropy for the quantum Ising and Heisenberg models on any hyperbolic ($p,r$) lattices AGquantum . The AdS/CFT correspondence in the current perspective lies in the idea of holographic entanglement entropy hp1 ; hp2 ; TNAds3 , where a non-gravitational theory can exist on the lattice boundary of hyperbolic $(p,r)$ lattices. The entanglement entropy originates in the reduced density matrix of the bulk subsystem measuring the amount of information within the AdS-CFT correspondence.
VI Acknowledgements
@@ D.1 Low-temperature limit
$\begin{split}{\cal F}_{4,\infty}&=\lim\limits_{k\to\infty}-\frac{k_{\rm B}T}{{\cal N}_{4,k}}\ln{\cal Z}_{4,k}\\ &=\lim\limits_{k\to\infty}-\frac{k_{\rm B}T}{{\cal N}_{4,k}}\ln\sum\limits_{[\sigma]}\exp\left(-\frac{{\cal H}_{p}[\sigma]}{k_{\rm B}T}\right)\\ &=\lim\limits_{k\to\infty}-\frac{J{\cal B}_{4,k}^{\leavevmode\nobreak\ }+|h|\,{\cal N}_{4,k}^{\leavevmode\nobreak\ }+|b|\,n_{4,k}^{\rm BS}}{{\cal N}_{4,k}}\\ &=\lim\limits_{k\to\infty}-\frac{J4k(2k-1)+|h|(2k)^{2}+|b|4(2k-1)}{(2k)^{2}}\\ &=-2J-|h|.\end{split}$ (43)
The boundary field $b$ does not affect the free energy in the thermodynamic limit because $n_{4,k}^{\rm BS}$ grows linearly, and $b$ is irrelevant on the Euclidean lattices.
Substantially different behavior occurs on the hyperbolic lattices ($p\geq 5$). For the Bethe lattice $(\infty,4)$ we obtain
$\begin{split}{\cal F}_{\infty,\infty}&=\lim\limits_{\begin{subarray}{c}k\to\infty\\ p\to\infty\end{subarray}}-\frac{k_{\rm B}T}{{\cal N}_{p,k}}\ln{\cal Z}_{p,k}\\ &=\lim\limits_{\begin{subarray}{c}k\to\infty\\ p\to\infty\end{subarray}}-\frac{J{\cal B}_{p,k}^{\leavevmode\nobreak\ }+|h|\,{\cal N}_{p,k}^{\leavevmode\nobreak\ }+|b|\,n_{p,k}^{\rm BS}}{{\cal N}_{p,k}}\\ &=\lim\limits_{\begin{subarray}{c}k\to\infty\\ p\to\infty\end{subarray}}-\frac{J(2p)^{k}+|h|(2p)^{k}+|b|(2p)^{k}}{(2p)^{k}}\\ &=-J-|h|-|b|.\end{split}$ (44)

### v2

@@ IV Results
We selected the clock and $q$-state Potts models because they exhibit first-order (discontinuous), second-order (continuous), and infinite-order Berezinskii-Kosterlitz-Thouless (BKT) BKT1 ; BKT2 ; BKT3 transitions on the square lattice $(4,4)$. However, the hyperbolic geometry affects the above results since the Hausdorff dimension of the hyperbolic lattices is infinite. To verify the vertex representation of TN on the hyperbolic lattices, we classify the spin models by phase transition, as stated by the weight representation Ueda1 ; weightpq ; Baxter . We proceed with the following:
@@
Now we check if the vertex representation of TN is correctly constructed. We calculate the phase-transition temperature of the Ising model for the Bethe lattice. The Bethe lattice with the coordination number $r=4$ can be accessed via the set of hyperbolic lattices and is equivalent to $(\infty,4)$ as also examined in Refs. weightpq, ; MFU, ; triangular, . The phase-transition temperature is exactly known Baxter to be $T_{\rm pt}^{(\infty,4)}=2/\ln(2)$. The convergence to the Bethe lattice is common for both the quantum and the classical spin models we have studied in our earlier works by the weight representation only AGquantum ; Mi1 ; Mi2 ; weightp4 .
Figure 8: The temperature dependence of the spontaneous magnetization towards the asymptotic Bethe lattice geometry $(\infty,4)$ with the bond dimension $m=4$. The left inset shows the fast asymptotic convergence of the entanglement entropy to the Bethe lattice. The right inset demonstrates the linearity, ${\cal M}^{2}_{p}\propto T$ leading to the mean-field exponent $\beta=\frac{1}{2}$.
@@
Therefore, we show Fig. 9 to display the entanglement entropy in the semi-logarithmic scale to emphasize the differences among the various $(p,4)$ lattices. The decreasing peaks of ${\cal E}_{p}$ completely coincide with the transition temperatures in ${\cal M}_{p}$. In both cases, ${\cal M}_{p}$ and ${\cal E}_{p}$ approach the exact value $T_{\rm pt}^{(\infty,4)}=2/\ln(2)$ equally rapidly on the Bethe lattice.
Figure 9: The entanglement entropy versus temperature in lin-log scale for lattices $(p,4)$ where $4\leq p\leq 50$ with the bond dimension $m=4$. The asymptotic convergence of $T_{\rm pt}^{(p,4)}$ to the Bethe lattice is well visible on the sharp maxima of ${\cal E}_{p}$.
Figure 10: Ising model ($q=2$ clock model) calculated on the hyperbolic $(20,4)$ lattice is numerically indistinguishable from the Bethe lattice. If fitting the mean-field exponent, we obtain $\beta=0.00002$ as can be read-off from the linear decrease of ${\cal M}_{20}^{2}$. The data points (asterisks) shown in the inset required $k>10^{11}$ iteration cycles to converge at bond dimension $m=2$.
@@ IV.3 Multi-state spin models
Figure 11: The entanglement entropy ${\cal E}_{5}$ vs temperature for the $q$-state clock models on the lattice $(5,4)$ with the bond dimension $m=q^{2}$. The inset shows the temperature dependence of spontaneous magnetization $M_{5}$ for the $q$-state clock models (lines) that coincides with the results of the weight-type TN, shown by the symbols ($\ast$).
We point out the non-diverging entropy ${\cal E}_{5}$ in the clock model for $q\geq 5$, as plotted in Fig. 11. Recall that ${\cal E}_{4}$ for the $q$-state clock models with $q\geq 2$ on the Euclidean ($4,4$) lattices logarithmically diverges at phase transitions MP ; Ueda ; CentrCharge . On the other hand, phase transitions on the hyperbolic lattices are always non-critical MFU ; triangular since the correlation length $\xi$ does not diverge Iharagi , reaches a maximal sharp peak and is small ($\xi<1$).
Entanglement entropy at phase transition for $h=b=0$ diverges logarithmically (${\cal E}_{4}\propto\frac{1}{12}\ln k$) on the Euclidean two-dimensional lattices CentrCharge , while ${\cal E}_{p\geq 5}$ is always finite on the hyperbolic lattices $0<{\cal E}_{{p\geq 5}}\ll 1$. The correlations decay exponentially if measured between the bulk and the lattice boundary. However, the correlations on the lattice boundary always decay as a power law, and the system is critical on the hyperbolic boundary TNAds1 . Comparing the qualitative similarity between the non-diverging ${\cal E}_{5}$ with the logarithmically diverging ${\cal E}_{4}$ on the square lattice MP , we can conjecture a non-critical BKT-like phase that exists in bulk only. However, no BKT phase transition is present for $p\geq 5$.
@@
Figure 12: The entanglement entropy ${\cal E}_{5}$ vs temperature for the $q$-state Potts models on $(5,4)$ lattice with the bond dimension $m=q^{2}$. We emphasize the emergence of the first-order transition for $q\geq 3$ on the hyperbolic lattices. The circles denote the phase-transition temperatures where the curves become discontinuous. Inset shows the equivalence of spontaneous magnetization ${\cal M}_{5}$ between the vertex TNs (lines) and the weight TNs ($\ast$).
@@ IV.4 Free Energy
Figure 13: The temperature dependence of the free energy for the Ising model ($q=2$) on $(p,4)$ lattices when imposing the boundary field $b=0$ and $b=\pm 2$ (at $h=0$). Calculated for the bond dimension $m=q^{2}$. The free energy does not depend on the boundary field $b$ on the square ($4,4$) lattice. The upper inset depicts a strictly linear dependence of the free energy on $b$ and $h$ at $T\to 0$ if studied on the Bethe lattice. The lower inset confirms the high-temperature regime of the thermodynamic entropy $S_{\rm th}\to\ln(q)$. The full and dashed lines correspond to the $q$-state clock and Potts models, respectively.
@@ V Conclusion
If comparing the current results via the vertex representation of the TN with the earlier weight representation of TN weightp4 ; weightpq , we confirm the full agreement between the two representations. Moreover, our results for the Ising model on the (5,4) lattice also agree with Monte Carlo (MC) simulations on a $(5,5)$ hyperbolic lattice QMC , where the authors also confirm the mean-field universality specifying phase-transition temperature.
The numerical accuracy using the CTMRG method prevails over the MC simulations for 2D lattices. While the outermost spin layers must be removed using MC, the CTMRG algorithm neglects boundary effects describing the bulk properties. Moreover, the authors QMC numerically calculated the mean-field universality exponent $\beta=0.46\pm 0.5$, whereas we achieved the exponent $\beta=0.500002$ almost reaching the exact value $\beta=\tfrac{1}{2}$. The deviations in MC arose due to finite-size scaling and the necessity to subtract a couple of boundary layers which significantly affect the bulk properties. In contrast, the most relevant errors in CTMRG originate from the round-off errors due to the large number of iterations (since the hyperbolic systems are off-critical and setting the bond dimensions $m\approx q^{2}$ is sufficient due to the exponential decay of density matrix eigenvalues Vvsweight ).
Having compared the most similar hyperbolic lattices (5, 5) calculated by MC and (5, 4) by CTMRG, we see that MC yields the phase-ransition temperature $T_{c}/J=3.93\pm 0.03$ on the (5, 5) lattice while CTMRG results in the phase transition $T_{c}/J=2.799083$ for (5, 4) lattice. The phase-transition temperature by CTMRG can be further improved if more iterations are taken, as demonstrated on the Bethe lattice in Fig. 10. Away from the phase transition, the numerical accuracy of all physical quantities reaches the machine precision. On the other hand, the MC simulations have been successfully applied to in the 3D hyperbolic space, and we hope to make improvements in the future). We notice that CTMRG is not accurate for the 3D cubic lattice where substantially large bond dimensions are necessary 3DCTMRG .
@@
The number of the physical spin indices $x_{\cal V}^{\leavevmode\nobreak\ }$, $x_{\cal P}^{\leavevmode\nobreak\ }$, $x_{\cal C}^{\leavevmode\nobreak\ }$ depends on the TN structure that is $p$, $2p$, $p(2p-7)$, respectively, in accord with the right-hand side of Fig. 15. Such a construction is inevitable for the calculation of $\langle\psi|{\cal H}|\psi\rangle_{(p,4)}$ and other observables on hyperbolic lattices.
VI Acknowledgments
@@ D.1 Low-temperature limit
$\begin{split}{\cal F}_{4,\infty}&=\lim\limits_{k\to\infty}-\frac{k_{\rm B}T}{{\cal N}_{4,k}}\ln{\cal Z}_{4,k}\\ &=\lim\limits_{k\to\infty}-\frac{k_{\rm B}T}{{\cal N}_{4,k}}\ln\sum\limits_{[\sigma]}\exp\left(-{{\cal H}_{p}[\sigma]}/{k_{\rm B}T}\right)\\ &=\lim\limits_{k\to\infty}-\frac{J{\cal B}_{4,k}^{\leavevmode\nobreak\ }+|h|\,{\cal N}_{4,k}^{\leavevmode\nobreak\ }+|b|\,n_{4,k}^{\rm BS}}{{\cal N}_{4,k}}\\ &=-2J-|h|.\end{split}$ (45)
The boundary field $b$ does not affect the free energy in the thermodynamic limit because $n_{4,k}^{\rm BS}$ grows linearly and $b$ is irrelevant on the Euclidean lattices.
Substantially different behavior occurs on the hyperbolic lattices. For the Bethe lattice $(\infty,4)$ we obtain
$\begin{split}{\cal F}_{\infty,\infty}&=\lim\limits_{\begin{subarray}{c}k\to\infty\\ p\to\infty\end{subarray}}-\frac{k_{\rm B}T}{{\cal N}_{p,k}}\ln{\cal Z}_{p,k}\\ &=\lim\limits_{\begin{subarray}{c}k\to\infty\\ p\to\infty\end{subarray}}-\frac{J{\cal B}_{p,k}^{\leavevmode\nobreak\ }+|h|\,{\cal N}_{p,k}^{\leavevmode\nobreak\ }+|b|\,n_{p,k}^{\rm BS}}{{\cal N}_{p,k}}\\ &=-J-|h|-|b|.\end{split}$ (46)